Probability Distribution Problem

1.) I throw a die and get P1.00 if it is showing 1, get P2.00 if it is showing 2, P3.00 if it

is showing 3, P4.00 if it is showing 4, P5.00 if it is showing 5, and get P6.00 if it is

showing 6. What is the amount of money I can expect if I throw it 100 times?

complete solution


1
Expert's answer
2021-05-26T03:34:51-0400

For one throw:

"E(X)=\frac{1}{6}(1+2+3+4+5+6)=\frac{21}{6}=P3.5."

For 100 throws:

"E(X)=3.5*100=P350."